A 140-g baseball traveling at 84 mi/h strikes the catchers mitt which, in bringing the ball to rest, recoils backward a distance of 20 cm. Determine the magnitude of the average force applied by the glove on the ball to bring it to a stop, in newtons. Answer: Check
Answer #1

Velocity, v = 84 mi/h = 37.55 m/s

Acceleration a = v^2/2s = 37.55^2/2 x 0.2 = 3525 m/s^2

Force, F = ma = 0.14 x 3525 = 493.50 N

Comment in case any doubt please rate my answer.....